Observe the following RXN:

CH4 + O2 --> CO + 2H2O

Carbon has a mass of 12 amu, Hydrogen 1 amu and Oxygen 16 amu.

If 16g of CH4 and 32g of O2 are reacted, how much CO is produced?
1
Expert's answer
2012-09-25T11:28:58-0400
The amount of CH4 is
n(CH4) = m(CH4)/M(CH4) = 16/16 = 1 mol
The amount of
O2 is
n(O2) = M(O2)/M(O2) = 32/32 = 1 mol
According to the reaction the
amount of CO is equal to the amount of CH4 and O2. Thus the amount of CO
is
n(CO) = n(O2) = n(CH4) = 1 mol
The mass of CO is
m(CO) = n(CO)*M(CO)
= 1*28 = 28 g